A half if what John's age was 4 years ago is equal to one third of what it will be in 5 years' time. How old is John now?
Answer by lwsshak3(6485)

A half if what John's age was 4 years ago is equal to one third of what it will be in 5 years' time. How old is John now?
**
let x= John's present age
x-4=John's age 4 years ago
x+5=John's age 5 years hence
..
(x-4)/2=(x+5)/3
3x-12=2x+10
x=22
ans:
John's present age=22